Use the given degree of confidence and sample data to construct a confidence interval for the population mean u. Assume that the poplution has a normal distribution
n=10 x=8.7, s=3.3, 95%confidence
1
Expert's answer
2011-04-05T10:41:18-0400
For normal distribution 95% interval Z = 1.96:

Upper limit:& x + 1.96*s/sqrt(n) = 8.7 + 1.96*3.3/sqrt(10) =10.75
Lower limit:& x - 1.96*s/sqrt(n) = 8.7 - 1.96*3.3/sqrt(10) =6.65